# Benchmarking LLMs via Uncertainty Quantification

### v2

@@ Abstract
The proliferation of open-source Large Language Models (LLMs) from various institutions has highlighted the urgent need for comprehensive evaluation methods. However, current evaluation platforms, such as the widely recognized HuggingFace open LLM leaderboard, neglect a crucial aspect – uncertainty, which is vital for thoroughly assessing LLMs. To bridge this gap, we introduce a new benchmarking approach for LLMs that integrates uncertainty quantification. Our examination involves nine LLMs (LLM series) spanning five representative natural language processing tasks. Our findings reveal that: I) LLMs with higher accuracy may exhibit lower certainty; II) Larger-scale LLMs may display greater uncertainty compared to their smaller counterparts; and III) Instruction-finetuning tends to increase the uncertainty of LLMs. These results underscore the significance of incorporating uncertainty in the evaluation of LLMs. [^1] Our implementation available at https://github.com/smartyfh/LLM-Uncertainty-Bench
[…]
Benchmarking LLMs via Uncertainty Quantification
[…]
@@ 1 Introduction
Large Language Models (LLMs) have gained significant traction within both academia and industry, with numerous organizations and companies open-sourcing their versions of LLMs (Chang et al., 2023; Zhao et al., 2023; Kaddour et al., 2023; Yang et al., 2023a). LLMs are highly versatile, demonstrating capabilities in various tasks such as question answering, document summarization, dialogue systems, and machine translation (Ye et al., 2023a; Pang et al., 2024). Given the growing interest and advancements in LLMs, it is crucial to establish appropriate methods for evaluating their performance (Liang et al., 2022; Ye et al., 2023b; Chang et al., 2023). However, conducting a comprehensive evaluation of LLMs remains a challenging endeavor (Guo et al., 2023; Zheng et al., 2024).
Figure 1: An illustration of two LLMs accurately predicting the true answer (with option A possessing the highest probability), but showing different levels of uncertainty. Note that when both LLMs predict a wrong answer, they may also display different uncertainties.
@@
To address this challenge, several open leaderboards such as the popular HuggingFace open LLM leaderboard [^2] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, OpenCompass (Contributors, 2023), Chatbot Arena (Zheng et al., 2024), and FlagEval (BAAI, 2023) have emerged, providing a comparative analysis of LLM performance. Despite their usefulness, these leaderboards possess a significant limitation: They do not take into account the uncertainty of LLMs. For example, the HuggingFace open LLM leaderboard only utilizes accuracy as the evaluation metric. However, as demonstrated in Figure 1, two LLMs may achieve identical accuracy scores but exhibit different levels of uncertainty regarding the question. This is analogous to students taking exams of multiple-choice questions, where two students may select the same answer but actually possess distinct degrees of uncertainty or comprehension about the question. Consequently, it is necessary to incorporate uncertainty into the evaluation process to achieve a more comprehensive assessment of LLMs.
@@
In this paper, we propose the utilization of conformal prediction (Vovk et al., 2005; Angelopoulos and Bates, 2021) as the method to quantify uncertainty in LLMs. Compared to alternative methods such as Bayesian variational inference (Hu et al., 2023), conformal prediction offers multiple advantages including ease of implementation, high efficiency, distribution-free and model-agnostic, and a statistically rigorous estimation of uncertainty rather than a heuristic approximation. Hence, conformal prediction can serve as a practical and principled means for assessing the uncertainty of LLMs.
@@
Specifically, we benchmark nine open-source LLMs (LLM series) across five typical Natural Language Processing (NLP) tasks, namely question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. Given that most existing open leaderboards and benchmarking datasets (Liu et al., 2024) focus on multiple-choice tasks, we also adopt the multiple-choice question setting for all tasks [^3] We leave the benchmarking of LLM uncertainty in the free-form text generation setting as our future work.. Although some of these tasks (e.g., document summarization) are inherently generative, it is challenging to develop a deterministic and reproducible method for quantifying the uncertainty within the generated text due to randomness in the generation process. Instead, we convert all tasks into multiple-choice questions, with the objective of each task being to select the correct option from the provided choices. Our empirical results reveal the following observations:
  • LLMs demonstrating higher accuracy may exhibit lower certainty.
  • LLMs with larger scales may display greater uncertainty than their smaller counterparts.
  • LLMs after instruction-finetuning tend to possess higher uncertainty.
@@ 2 Related Work
Uncertainty quantification (Fomicheva et al., 2020; Abdar et al., 2021; Gawlikowski et al., 2023) has been an active area of research in both machine learning and NLP due to its importance in real-world applications such as decision making, risk assessment, human-AI collaboration, and so on. Typical uncertainty quantification methods include confidence-based methods (Hu et al., 2023), Bayesian methods (Kwon et al., 2020), and ensemble methods (Rahaman et al., 2021). Confidence-based methods such as entropy can be sensitive to poor calibration and may not fully capture models’ underlying uncertainties (Minderer et al., 2021). Bayesian methods and ensemble methods usually suffer from high computational complexity (Gawlikowski et al., 2023), making them not suitable for assessing the uncertainty of LLMs.
@@
Recently, there has been a growing interest in applying conformal prediction for uncertainty quantification (Angelopoulos and Bates, 2021; Kumar et al., 2023; Quach et al., 2023; Lu et al., 2023). For example, conformal prediction has been applied to part-of-speech prediction (Dey et al., 2022), paraphrase detection (Giovannotti and Gammerman, 2021), and fact verification (Fisch et al., 2020). Similar to the process of elimination used by students during exams, conformal prediction identifies a subset of potential labels in classification tasks by excluding improbable labels, which is statistically guaranteed to contain the true label, and quantifies uncertainty as the size of this subset (Angelopoulos et al., 2020). The coverage guarantee property makes conformal prediction a highly robust uncertainty quantification method. In addition, conformal prediction is non-parametric, distribution-free, model-agnostic, and computationally efficient. Therefore, it is a favorable choice in the context of LLMs.
@@
Evaluating the performance of LLMs is a crucial aspect of their development and deployment (Huang et al., 2023). Current studies assess LLMs from different angles using specific datasets, such as MMLU (Hendrycks et al., 2020) for knowledge, HellaSwag (Zellers et al., 2019) for reasoning, HaluEval Li et al. (2023a) for hallucination, GSM8K (Cobbe et al., 2021) for math, and BOLD (Dhamala et al., 2021) for fairness. Besides, evaluation platforms like HuggingFace open LLM leaderboard and Chatbot Arena (Zheng et al., 2024) have also been developed to facilitate comparisons among LLMs. Despite these efforts, the critical aspect of uncertainty in LLMs remains underexplored. More recently, some research has begun to consider uncertainty in LLMs (Xiong et al., 2023; Yang et al., 2023b; Lin et al., 2023; Chen and Mueller, 2023; Wagle et al., 2023; Fadeeva et al., 2023). However, these approaches such as the sampling-based semantic entropy (Kuhn et al., 2023) are heuristic and lack a standardized methodology for benchmarking purposes. In contrast, our utilization of conformal prediction can provide a robust and systematic evaluation of uncertainty.
@@ 3 Background of Conformal Prediction
Conformal prediction serves as a distribution-free and model-agnostic approach to uncertainty quantification (Vovk et al., 2005; Balasubramanian et al., 2014; Angelopoulos and Bates, 2021; Fontana et al., 2023). It can transform any heuristic notion of uncertainty from any model into a statistically rigorous one. As aforementioned, for multi-class classification tasks, conformal prediction outputs a prediction set of possible labels (answers) that encompasses the correct label with a user-specified error rate and expresses uncertainty as the set size. Intuitively, a larger set size indicates higher uncertainty and vice versa.
@@
Equation (1) requires that the generated prediction set should contain the true label $Y_{t}$ with a probability no smaller than $1-\alpha$. This coverage guarantee requirement can be achieved with the aid of a small amount of held-out calibration data $\mathcal{D}_{cal}=\{(X^{(1)}_{c},Y^{(1)}_{c}),\dots,(X^{(n)}_{c},Y^{(n)}_{c})\}$, where $n$ denotes the number of data points in the calibration set [^4] It is also required that the test data points are drawn from the same distribution as the calibration data.. More specifically, conformal prediction works in the following process (Angelopoulos and Bates, 2021) to create the prediction set:
@@
LAC (Sadinle et al., 2019) defines the conformal score function as
@@
where $f(X)_{Y}$ is the softmax score corresponding to the true label. It has been proven that LAC can lead to prediction sets with the smallest average size (Sadinle et al., 2019). However, it may undercover hard instances and overcover easy ones.
@@
APS (Romano et al., 2020) defines the conformal score function as
@@ 4 Evaluation Tasks and Datasets
LLMs have demonstrated remarkable capabilities across various aspects (Guo et al., 2023; Naveed et al., 2023). It is essential to develop multiple tasks to evaluate their performance comprehensively. For this purpose, we consider five typical NLP tasks, including question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. For each task, we prepare a dataset with 10,000 instances. In addition, we formulate each task as a multiple-choice question answering (MCQA) task and the objective is to select the only correct answer out of six possible options (i.e. A, B, C, D, E, and F). It is worth emphasizing that the prevailing benchmarking open leaderboards and datasets also focuses on MCQA tasks (Guo et al., 2023; Liu et al., 2024).
@@
QA is applied to evaluate an LLM’s proficiency in utilizing its extensive world knowledge to provide accurate answers to a diverse range of questions. For this task, we adopt MMLU (Hendrycks et al., 2020) as the evaluation dataset. MMLU encompasses a total of 57 subjects, spanning various disciplines such as elementary mathematics, US history, computer science, and law. These subjects are further classified into four broad categories, namely humanities, social sciences, STEM, and others (business, health, misc.). For each category, we sample 2500 instances, leading to 10,000 instances in total.
@@
RC is used for testing an LLM’s ability to understand and analyze a given context, comprehend the meaning of words and sentences, and answer questions based on the information presented in the context. It also tests the ability of LLMs to make inferences and draw conclusions from the given context. We take CosmosQA (Huang et al., 2019) as the evaluation dataset. CosmosQA focuses on reading between the lines (Norvig, 1987) over a diverse collection of people’s everyday narratives that require reasoning beyond the exact text spans in the context. Due to the unavailability of ground truth labels for the test set in CosmosQA, we sample 10,000 instances from the training and development sets.
@@
CI is leveraged to evaluate the ability of LLMs to understand and reason about the relationships between concepts and events based on commonsense and background knowledge. This type of testing helps to assess the generalization and reasoning abilities of LLMs beyond simple pattern recognition tasks. We employ HellaSwag (Zellers et al., 2019) as the evaluation dataset. HellaSwag focuses on commonsense natural language inference whose target is to select the most likely followup to a given event description. Same as CosmosQA, we sample 10,000 instances from the training and development sets of HellaSwag for the purpose of evaluation.
@@
DRS is adopted for assessing the ability of LLMs to comprehend the meaning of a given dialogue and select an appropriate response from a set of possible responses. This includes the ability to comprehend the meaning of the user’s input, select appropriate responses that are relevant to the conversational context, and maintain coherence and consistency in the dialogue. We utilize the dialogue data from the HaluEval Li et al. (2023a) benchmark as the evaluation dataset (denoted as HaluDial). This dataset consists of exactly 10,000 instances and is built upon OpenDialKG (Moon et al., 2019), a knowledge-grounded dialogue dataset.
@@
DS is taken to evaluate the proficiency of LLMs in comprehending the substance and context of a given document, and in producing a succinct and cohesive summary that effectively conveys the crucial information and main ideas of the document. This requires the LLM to have a good understanding of the language, the topic, and the structure of the document. Similar to DRS, we adopt the summarization data from the HaluEval Li et al. (2023a) benchmark as the evaluation dataset (denoted as HaluSum). This dataset also comprises precisely 10,000 instances and is derived from CNN/Daily Mail (See et al., 2017), a summarization dataset pertaining to news articles.
@@ 5 Evaluation Prompts and Metrics
As delineated in Section 3, one crucial step of leveraging conformal prediction for uncertainty quantification is to acquire the softmax score corresponding to each option. Next, we explicate the method used to elicit these scores from LLMs, followed by a description of the evaluation metrics utilized.
5.1 Prompting Strategies
Following previous works (Hendrycks et al., 2020; Gao et al., 2023; Zhang et al., 2023; Chen et al., 2023; Zheng et al., 2024), we rely on prompt engineering rather than supervised finetuning as the testing approach to evaluating the performance of LLMs on each task. However, our preliminary results show that LLMs are sensitive to prompts. In this regard, we consider three prompting strategies, including Base Prompt, Shared Instruction Prompt, and Task-specific Instruction Prompt, to reduce the influence of LLMs’ sensitivity to different prompts, thereby ensuring a fairer comparison.
@@ 5.1 Prompting Strategies
5.2 Evaluation Metrics
We evaluate LLMs from two perspectives, namely prediction accuracy and prediction uncertainty. For prediction accuracy, we adopt the commonly used metric – Accuracy (Acc). To evaluate prediction uncertainty, we use Set Size (SS), which is a primary metric for conformal prediction (Angelopoulos and Bates, 2021). Let $Y_{p}$ be the prediction for the test instance $(X_{t},Y_{t})\in\mathcal{D}_{test}$. These two metrics can be calculated as follows:
$Acc=\frac{1}{|\mathcal{D}_{test}|}\sum_{(X_{t},Y_{t})\in\mathcal{D}_{test}}\mathbbm{1}(Y_{p}=Y_{t}),$ (6)
$SS=\frac{1}{|\mathcal{D}_{test}|}\sum_{(X_{t},Y_{t})\in\mathcal{D}_{test}}|\mathcal{C}(X_{t})|,$ (7)

### v3

@@ Abstract
The proliferation of open-source Large Language Models (LLMs) from various institutions has highlighted the urgent need for comprehensive evaluation methods. However, current evaluation platforms, such as the widely recognized HuggingFace open LLM leaderboard, neglect a crucial aspect – uncertainty, which is vital for thoroughly assessing LLMs. To bridge this gap, we introduce a new benchmarking approach for LLMs that integrates uncertainty quantification. Our examination involves nine LLMs (LLM series) spanning five representative natural language processing tasks. Our findings reveal that: I) LLMs with higher accuracy may exhibit lower certainty; II) Larger-scale LLMs may display greater uncertainty compared to their smaller counterparts; and III) Instruction-finetuning tends to increase the uncertainty of LLMs. These results underscore the significance of incorporating uncertainty into the evaluation of LLMs. Our implementation is available at https://github.com/smartyfh/LLM-Uncertainty-Bench.
@@ 1 Introduction
Large Language Models (LLMs) have gained significant traction within both academia and industry, with numerous organizations and companies open-sourcing their versions of LLMs [9, 74, 36, 68]. LLMs are highly versatile, demonstrating capabilities in various tasks such as question answering, document summarization, dialogue systems, and machine translation [70, 52]. Given the growing interest and advancements in LLMs, it is crucial to establish appropriate methods for evaluating their performance [42, 71, 9]. However, conducting a comprehensive evaluation of LLMs remains a challenging endeavor [28, 75].
@@
To address this challenge, several open leaderboards such as the popular HuggingFace open LLM leaderboard, [^1] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard OpenCompass [14], Chatbot Arena [75], and FlagEval [6] have emerged, providing a comparative analysis of LLM performance. Despite their usefulness, these leaderboards possess a significant limitation: They do not take into account the uncertainty of LLMs. For example, the HuggingFace open LLM leaderboard only utilizes accuracy as the evaluation metric. However, as demonstrated in Figure 1, two LLMs may achieve identical accuracy scores but exhibit different levels of uncertainty regarding the question. This is analogous to students taking exams of multiple-choice questions, where two students may select the same answer but actually possess distinct degrees of uncertainty or comprehension about the question. Consequently, it is necessary to incorporate uncertainty into the evaluation process to achieve a more comprehensive assessment of LLMs.
@@
In this paper, we propose the utilization of conformal prediction [64, 5] as the method to quantify uncertainty in LLMs. Compared to alternative methods such as Bayesian variational inference [30], conformal prediction offers multiple advantages including ease of implementation, high efficiency, distribution-free and model-agnostic, and a statistically rigorous estimation of uncertainty rather than a heuristic approximation [5]. Hence, conformal prediction can serve as a practical and principled means for assessing the uncertainty of LLMs.
Figure 1: An illustration of two LLMs accurately predicting the true answer (with option A possessing the highest probability), but showing different levels of uncertainty. Note that when both LLMs predict a wrong answer, they may also display different levels of uncertainty.
@@
Specifically, we benchmark nine open-source LLMs (LLM series) across five typical Natural Language Processing (NLP) tasks, namely question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. Given that most existing open leaderboards and benchmarking datasets [45] focus on multiple-choice tasks, we also adopt the multiple-choice question setting for all tasks. Although some of these tasks (e.g., document summarization) are inherently generative, it is challenging to develop a deterministic and reproducible method for quantifying the uncertainty within the generated text due to randomness in the generation process. Instead, we convert all tasks into multiple-choice questions, with the objective of each task being to select the correct option from the provided choices. Our empirical results reveal the following observations: I) LLMs demonstrating higher accuracy may exhibit lower certainty; II) LLMs with larger scales may display greater uncertainty than their smaller counterparts; and III) LLMs after instruction-finetuning tend to possess higher uncertainty.
@@ 2 Related work
Uncertainty quantification [22, 2, 25] has been an active area of research in both machine learning and NLP due to its importance in real-world applications such as decision making, risk assessment, human-AI collaboration, and so on. Typical uncertainty quantification methods include confidence-based methods [30], Bayesian methods [39], and ensemble methods [54]. Confidence-based methods such as entropy can be sensitive to poor calibration and may not fully capture models’ underlying uncertainties [48]. Bayesian methods and ensemble methods usually suffer from high computational complexity [25], making them not suitable for assessing the uncertainty of LLMs.
@@
Recently, there has been a growing interest in applying conformal prediction for uncertainty quantification [5, 38, 53, 46]. For example, conformal prediction has been applied to part-of-speech prediction [17], paraphrase detection [26], and fact verification [21]. Similar to the process of elimination used by students during exams, conformal prediction identifies a subset of potential labels in classification tasks by excluding improbable labels, which is statistically guaranteed to contain the true label, and quantifies uncertainty as the size of this subset [4]. The coverage guarantee property makes conformal prediction a highly robust uncertainty quantification method. In addition, conformal prediction is non-parametric, distribution-free (i.e. not dependent on any specific distributional assumptions about the data), model-agnostic, and computationally efficient [5]. Therefore, it is a favorable choice in the context of LLMs.
@@
Evaluating the performance of LLMs is a crucial aspect of their development and deployment [32]. Current studies assess LLMs from different angles using specific datasets, such as MMLU [29] for knowledge, HellaSwag [72] for reasoning, HaluEval [40] for hallucination, GSM8K [13] for math, and BOLD [18] for fairness. Besides, evaluation platforms like HuggingFace open LLM leaderboard and Chatbot Arena [75] have also been developed to facilitate comparisons among LLMs. Despite these efforts, the critical aspect of uncertainty in LLMs remains underexplored. More recently, some research has begun to consider uncertainty in LLMs [67, 69, 43, 10, 65, 20]. However, these approaches such as the sampling-based semantic entropy [37] are heuristic and lack a standardized methodology for benchmarking purposes. In contrast, our utilization of conformal prediction can provide a robust and systematic evaluation of uncertainty.
@@ 3 Background of conformal prediction
Conformal prediction serves as a distribution-free and model-agnostic approach to uncertainty quantification [64, 8, 5, 23]. It can transform any heuristic notion of uncertainty from any model into a statistically rigorous one. As aforementioned, for multi-class classification tasks, conformal prediction outputs a prediction set of possible labels (answers) that encompasses the correct label with a user-specified error rate and expresses uncertainty as the set size. Intuitively, a larger set size indicates higher uncertainty and vice versa.
@@
Equation (1) requires that the generated prediction set should contain the true label $Y_{t}$ with a probability no smaller than $1-\alpha$. This coverage guarantee requirement can be achieved with the aid of a small amount of held-out calibration data $\mathcal{D}_{cal}=\{(X^{(1)}_{c},Y^{(1)}_{c}),\dots,(X^{(n)}_{c},Y^{(n)}_{c})\}$, where $n$ denotes the number of data points in the calibration set. [^2] It is also required that the test data points are drawn from the same distribution as the calibration data. More specifically, conformal prediction works in the following process [5] to create the prediction set:
@@
LAC [58] defines the conformal score function as
@@
where $f(X)_{Y}$ is the softmax score corresponding to the true label. It has been proven that LAC can lead to prediction sets with the smallest average size [58]. However, it may undercover hard instances and overcover easy ones.
@@
APS [57] defines the conformal score function as
@@ 4 Evaluation tasks and datasets
LLMs have demonstrated remarkable capabilities across various aspects [28, 50]. It is essential to develop multiple tasks to evaluate their performance comprehensively. For this purpose, we consider five typical NLP tasks, including question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization. For each task, we prepare a dataset with 10,000 instances. In addition, we formulate each task as a Multiple-Choice Question Answering (MCQA) task and the objective is to select the only correct answer out of six possible options (i.e. A, B, C, D, E, and F). It is worth emphasizing that the prevailing benchmarking open leaderboards and datasets also focus on MCQA tasks [28, 45].
@@
QA is applied to evaluate an LLM’s proficiency in utilizing its extensive world knowledge to provide answers to a diverse range of questions. For this task, we adopt MMLU [29] as the evaluation dataset. MMLU encompasses a total of 57 subjects, spanning various disciplines such as elementary mathematics, US history, computer science, and law. These subjects are further classified into four broad categories, namely humanities, social sciences, STEM, and others (business, health, misc.). For each category, we sample 2500 instances, leading to 10,000 instances in total.
@@
RC is used for testing an LLM’s ability to understand and analyze a given context, comprehend the meaning of words and sentences, and answer questions based on the information presented in the context. It also tests the ability of LLMs to make inferences and draw conclusions from the given context. We take CosmosQA [31] as the evaluation dataset. CosmosQA focuses on reading between the lines [51] over a diverse collection of people’s everyday narratives that require reasoning beyond the exact text spans in the context. Due to the unavailability of ground truth labels for the test set in CosmosQA, we sample 10,000 instances from the training and development sets.
@@
CI is leveraged to evaluate the ability of LLMs to understand and reason about the relationships between concepts and events based on commonsense and background knowledge. This type of testing helps to assess the generalization and reasoning abilities of LLMs beyond simple pattern recognition tasks. We employ HellaSwag [72] as the evaluation dataset. HellaSwag focuses on commonsense natural language inference whose target is to select the most likely followup to a given event description. Same as CosmosQA, we sample 10,000 instances from the training and development sets of HellaSwag for the purpose of evaluation.
@@
DRS is adopted for assessing the ability of LLMs to comprehend the meaning of a given dialogue and select an appropriate response from a set of possible responses. This includes the ability to comprehend the meaning of the user’s input, select appropriate responses that are relevant to the conversational context, and maintain coherence and consistency in the dialogue. We utilize the dialogue data from the HaluEval [40] benchmark as the evaluation dataset (denoted as HaluDial). This dataset consists of exactly 10,000 instances and is built upon OpenDialKG [49], a knowledge-grounded dialogue dataset.
@@
DS is taken to evaluate the proficiency of LLMs in comprehending the substance and context of a given document, and in producing a succinct and cohesive summary that effectively conveys the crucial information and main ideas of the document. This requires the LLM to have a good understanding of the language, the topic, and the structure of the document. Similar to DRS, we adopt the summarization data from the HaluEval [40] benchmark as the evaluation dataset (denoted as HaluSum). This dataset also comprises precisely 10,000 instances and is derived from CNN/Daily Mail [59], a summarization dataset pertaining to news articles.
@@ 5 Evaluation prompts and metrics
As delineated in § 3, one crucial step of leveraging conformal prediction for uncertainty quantification is to acquire the softmax score corresponding to each option. Next, we explicate the method used to elicit these scores from LLMs, followed by a description of the evaluation metrics utilized.
Prompting Strategies
Following previous works [29, 24, 73, 12, 75], we rely on prompt engineering rather than supervised finetuning as the testing approach to evaluating the performance of LLMs on each task. However, our preliminary results show that LLMs are sensitive to prompts. In this regard, we consider three prompting strategies, including Base Prompt, Shared Instruction Prompt, and Task-specific Instruction Prompt, to reduce the influence of LLMs’ sensitivity to different prompts, thereby ensuring a fairer comparison.
@@
Evaluation Metrics
We evaluate LLMs from two perspectives, namely prediction accuracy and prediction uncertainty. For prediction accuracy, we adopt the commonly used metric – Accuracy (Acc). To evaluate prediction uncertainty, we use Set Size (SS), which is a primary metric for conformal prediction [5]. Let $Y_{p}$ be the prediction for the test instance $(X_{t},Y_{t})\in\mathcal{D}_{test}$. These two metrics can be calculated as follows:
$Acc=\frac{1}{|\mathcal{D}_{test}|}\sum_{(X_{t},Y_{t})\in\mathcal{D}_{test}}\mathbbm{1}(Y_{p}=Y_{t}),\quad SS=\frac{1}{|\mathcal{D}_{test}|}\sum_{(X_{t},Y_{t})\in\mathcal{D}_{test}}|\mathcal{C}(X_{t})|,$ (6)